Molality & Molarity 1.


Calculate the molality of a solution containing 10.8g of ethylene glycol (C2H6O2) in 360g of water.

1
Expert's answer
2022-03-12T02:00:27-0500

Solution:

solute = ethylene glycol (C2H6O2)

solvent = water (H2O)


The molar mass of ethylene glycol (C2H6O2) is 62.07 g/mol

Therefore,

Moles of C2H6O2 = (10.8 g C2H6O2) × (1 mol C2H6O2 / 62.07 g C2H6O2) = 0.174 mol C2H6O2


Convert g to kg:

Kilograms of solvent = (360 g H2O) × (1 kg / 1000 g) = 0.36 kg H2O


Molality = Moles of solute / Kilograms of solvent

Therefore,

Molality of C2H6O2 solution = Moles of C2H6O2 / Kilograms of solvent

Molality of C2H6O2 solution = (0.174 mol) / (0.36 kg) = 0.483 mol/kg = 0.483 m

Molality of C2H6O2 solution = 0.483 m


Answer: The molality of C2H6O2 solution is 0.483 m